OpenStudy (anonymous):

The city of Atlanta recently increased their sales tax from 7.5% to 8%. How much more tax would you pay on a $10 purchase?

jimthompson5910 (jim_thompson5910):

The difference from 7.5% to 8% is 0.5% since 8 - 7.5 = 0.5 Take 0.5% of $10 to get 0.5% of $10 = (0.5/100)*10 = ??? tell me what you get

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

what is `0.5/100` equal to?

jimthompson5910 (jim_thompson5910):

I think you wrote something? but you deleted it? anyways, you should find that 0.5/100 = 0.005

jimthompson5910 (jim_thompson5910):

now just multiply that with 10 to get 0.005*10 = ???

OpenStudy (anonymous):

0.05?

jimthompson5910 (jim_thompson5910):

yep so overall, 0.5% of $10 is equal to $0.05 which is equivalent to 5 cents

jimthompson5910 (jim_thompson5910):

if the tax goes from 7.5% to 8%, then you pay an additional 5 cents